On 1/1/2020, SF acquired 25% of the outstanding common stock of LA, by purchasing 5,000 shares of the outstanding common stock of LA at a cost of $100,000. SF paid cash for the investment in LA on 1/1/2020. LA had a total of 20,000 shares of outstanding common stock. The total fair value of the 20,000 shares of LA's common stock on 1/1/2020, was $400,000. The total fair value of the 20,000 shares of LA's common stock on 12/31/2020 is $600,000. Both companies’ fiscal years end on 12/31. The following data pertains to LA during 2020.

Per share dividend declared and paid between 1/1/2020 and 12/31/2020: $2.00

Net income for 2020: 200,000

What is the balance (or value) of investment in LA in SF’s balance sheet as of 12/31/2020? Also, prepare all journal entries for SF in relation to the transactions of its investment in LA during 2020.
Answer #1

Answer :

The current invesment is valued at Fair value or Cost, whichever is lower.

Fari value at 31/12/2020 of 5000 shares = 600000/20000*5000 = 150000

Cost = 100000

Value to be shown in BS = 100000

Jounal Entries:

Date Debit credit 100000 100000 100000 Partculars 01-01-20 Investment in Shares A/C To Broker A/C (Being Shares purchsed from

Hope the above meet your objective

In case of any doubt reach us in comment section